OpenStudy (anonymous):

Factor completely p^4 – 18p^2 + 81

OpenStudy (anonymous):

\(p^4 – 18p^2 + 81=(p^2-9)^2=[(p-3)(p+3)]^2=(p-3)^2(p+3)^2.\)

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

You're welcome!